OpenStudy (anonymous):

Tom drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 9 hours. When Tom drove home, there was no traffic and the trip only took 4 hours. If his average rate was 40 miles per hour faster on the trip home, how far away does Tom live from the mountains? Do not do any rounding.

OpenStudy (wolf1728):

distance = rate * time distance = (r-40) * 9 distance = r * 4 distance = (r-40) * 9 distance = (9r-360) distance = r * 4 (9r-360) = r *4 5r -360 = 0 r = 72 mph Trip to = 32 mph * 9 hours = 288 miles Return Trip = 72 * 4 hours = 288 miles So distance = 288 miles Wow driving at 72 mph !! (That's somewhat over the extremely "fair" 55 mph speed limit.) LOL